# Superconducting resonator parametric amplifiers with intrinsic separation of pump and signal tones

### v2

@@ I Introduction
Signal amplification in superconducting parametric amplifiers is achieved through wave-mixing processes induced by the underlying nonlinearity, both reactive and dissipative, of the superconducting thin-films [1, 8, 9]. Experimentally, kinetic inductance amplifiers have demonstrated gains of $10-30\,\mathrm{dB}$ whilst achieving noise performance close to the standard quantum limit for linear amplifiers of half a quantum [10, 1, 11, 12]. By operating in degenerate mixing mode, these amplifiers have also been used to realise squeezed amplification and generate squeezed vacuum states [13]. Superconducting parametric amplifiers have been realised in both travelling-wave and resonator geometries. In general, amplifiers based on travelling-wave designs have wider bandwidths, whereas amplifiers based on resonator designs have lower pump power requirements and are less prone to lithographic defects such as shorts or breaks on long ($>0.5\,\mathrm{m}$) transmission lines with micron-scale features [1, 14, 2]. Regardless of geometry, central to the process of parametric amplification is the use of a strong pump tone: the nonlinear wave-mixing processes typically transfer energy from the strong pump tone to the weak signal tone in order to achieve amplification. It is important that this strong pump tone (as large as $-8.0\,\mathrm{dBm}$ at the input of the amplifiers in some instances [1]) is separated from the amplified signal tone so that no unwanted saturation or nonlinearity occurs in the downstream electronics, especially cryogenic HEMT amplifiers. Further, in the case of four-wave-mixing, the pump frequency is located close to the peak of the signal gain profile. As a result, this region of maximum gain may be unusable due to contamination of the signal by the phase-noise of the pump.
@@ II Theory
Figure 1: Top subfigure: schematic diagram of the pump separation scheme. The pump tone and the signal tone are injected into port 1 of the circulator; the input tones are sent to the half-wave resonator amplifier which is connected to port 2 of the circulator; the transmission mode of the resonator amplifier, which contain both the amplified signal tone and the pump tone, are terminated at a matched $50\,\mathrm{\Omega}$ termination; the reflection mode of the resonator amplifier, which contains mostly the amplified signal tone, is received via port 3 of the circulator as output. Bottom subfigure: an example of the half-wave resonators measured in this study. The resonator comprises a length of coplanar waveguide capacitively coupled in series between two contact pads. The length over which the coupling pad overlaps with the resonator line determines the coupling quality factor of this resonator.
@@
Figure 2: Squared S-parameter of the reflection mode of a half-wave two port resonator at different sweep powers, simulated using the theory presented in [20]. The sweep direction is from low frequency to high frequency, and the power increases from red to magenta. Accordingly, the resonance response shows increasing distortion due to a reactive feedback process. Discontinuity in the resonance response develops when the sweep power is greater than that required to induce bifurcation.
The reactive, dissipative, and rate-limited nonlinear response of a superconducting resonator to a strong pump has been well-described recently [18, 19]. In the presence of a strong sweeping tone, the stored and dissipated energies modify the circuit parameters of the resonator, which in turn change the frequency of the underlying, not directly observed, resonance. This theory has been used to describe accurately key features such as shifts in resonance frequencies, distortions in S-parameter profiles, and hysteresis and bifurcations: i.e. the resonator responds differently depending on whether the frequency is swept up or swept down. As shown in figure 2, as the sweep power is increased, from the red line to the magenta line, the resonance shifts and distorts largely as a result of reactive parametric feedback. When the applied power is greater than that required for bifurcation, a discontinuity appears in the observed transmission profile. As described by our recent work on resonator parametric amplifiers [18], the onset of this discontinuity is associated with the formation of operating points that can be used for high-gain, low-noise amplification. In practice, the optimum pump power is approximately the applied power at bifurcation, and the optimum pump frequency is approximately the frequency at which the discontinuity first appears.
@@
The amplification of a signal tone in a superconducting resonator has been previously described using an expanded nonlinear resonator model [20], which takes into account the effect of reactive, dissipative, and rate-limited nonlinearities in the wave-mixing process. Crucially, the behaviour of the weak signal tone is changed significantly by nonlinear wave-mixing between the signal and the pump. At high gain, this process generates amplified signals of similar magnitude propagating in both the transmission and the reflection directions. We have simulated the signal amplification and pump response using [20] at a fixed operating point, where we set the total quality factor $Q_{\mathrm{r}}$ to be $1000$ and coupling quality factor $Q_{\mathrm{c}}$ to be $1100$ such that the simulation models a realistic device fabricated in our laboratory. The top subfigure of figure 3 shows the resonance response in reflection where the sweep power is equal to the pump power of the amplifier’s operating point; the bottom subfigure shows the gain profile of the reflected signal tone as a function of frequency. The peak gain in the bottom subfigure is $17\,\mathrm{dB}$. The dashed line indicates the best operating frequency of the pump tone; in the top subfigure, it intersects the pump response at a frequency just above the discontinuity with a reflection coefficient of $-10.5\,\mathrm{dB}$. This simulation illustrates that by operating a half-wavelength reflection amplifier with pump frequency just above the bifurcation frequency, the pump power reflected is significantly weaker than the pump power incident on the device; the signal tone, however, is strongly amplified in both the transmission mode and the reflection mode.
@@
Finding the operating point of a superconducting resonator parametric amplifier appears, at first sight, to be a complicated task. As a strong pump is swept across the notional resonance frequency, the energies stored and dissipated in the resonator change. Although the underlying resonance simply shifts and broadens, retaining a simple Lorentzian form, this is not what is observed by the experimenter in a swept measurement, which exhibits distortion and bifurcation. This effect is highly reproducible, and well described by our theory [20], and so we have been able to establish a set-up procedure, which has proven to be exceedingly dependable and reproducible when applied to a wide range of resonator configurations and materials. The ruggedness of the operation is one of the key attractions of the resonator approach. The experimental procedure is as follows:
@@
III Devices and measurements
We have designed, fabricated, and packaged half-wave NbN resonators to experimentally realise the pump separation scheme. NbN films were deposited onto $50.8\,\,\mathrm{mm}$ diameter Si wafers by DC magnetron sputtering at a base pressure of $2\times 10^{-10}\,\,\mathrm{Torr}$ or below. NbN films were deposited by reactive sputtering a Nb target in the presence of $\mathrm{N_{2}}$ and the films were patterned using an image-reversal lift-off process. These NbN films undergo superconducting transition at temperature of $\sim 10\,\mathrm{K}$. All devices tested in this study were half-wave resonators based on co-planar waveguide (CPW) geometries, as shown in the bottom subfigure of figure 1. The CPW conductor width was $5\,\mathrm{\mu m}$, and the gap width from conductor to ground was $50\,\mathrm{\mu m}$. These resonators were end coupled through capacitive small re-entrant lines. The film thicknesses were controlled by calibrating and varying the deposition time. All films in this study had thicknesses of $\sim 100\,\mathrm{nm}$. The wafers were diced into individual resonators, which were bonded using Al wires to a gold-plated copper enclosure with SMA connectors. The enclosure was designed such that its resonance frequencies do not overlap significantly with the resonance frequencies of the half-wave resonators [21]. The packaged amplifier and the cryogenic circulator were arranged according to the top subfigure of figure 1. Both items were attached to the cold stage of an adiabatic demagnetization refrigerator. Stainless steel coaxial cables were used in the cryogenic system, thermally anchored at $35\,\mathrm{K}$, $4\,\mathrm{K}$, $1\,\mathrm{K}$, and $50\,\mathrm{mK}$ stages. HEMT amplifiers were not used in the signal chain since added-noise temperature measurement was not an aim of this study. The pump tone was generated and measured using a vector network analyser (VNA) and the signal tone was generated and measured using another VNA. The two VNAs were phase-locked to a common $10\,\mathrm{MHz}$ reference tone. The signal and the pump tone were combined at the input of the cryostat using a multiband two-way power combiner. Unless otherwise specified, all measurements were taken at cryogenic temperature of $0.1\,\mathrm{K}$, maintained using an adiabatic demagnetization refrigerator with a pulse tube cooler.
III.1 Proof-of-concept measurements
@@
In our first set of experiments, we applied the pump separation scheme to a half-wave NbN resonator with quality factor of $Q_{\mathrm{r}}\sim 30000$ (when measured using a weak sweep tone). Figure 4 shows measurements of S-parameter where the sweep tone was injected via port 1 of the circulator and measured via port 3 of the circulator. The ports were configured according to figure 1. The measured values were offset by a $20\,\mathrm{dB}$ attenuator before the input of the cryostat, coaxial cable loss in the cryostat, and circulator loss at the cold stage of the cryostat. (This is the only measurement where there is an attenuator in the pump path. The attenuator was included here to enable measurement of the resonance at very low sweep powers.) The shift in resonance frequency was initially in the positive frequency direction and then changed back to the negative frequency direction as the power was increased further. This complicated effect has previously been observed and discussed in [9] in the context of Ti, Nb, and NbN resonators. As seen in figure 4, when the sweep power increases to $2\,\mathrm{dBm}$ and above, the S-parameter shows a discontinuity, which is the reactive bifurcation. Importantly, we also notice that the depth of resonance becomes shallower as the sweep power increases. This indicates a decrease in the internal quality factor of the resonator as sweep power increases. From well-established resonator theories [19, 18], the depth of the resonance is given by
$S^{2}$ $=\left(1-\frac{Q_{\mathrm{r}}}{Q_{\mathrm{c}}}\right)^{2}\,,$ (1)
where $Q_{\mathrm{r}}$ is the total quality factor given by $Q_{\mathrm{r}}^{-1}=Q_{\mathrm{i}}^{-1}+Q_{\mathrm{c}}^{-1}$, $Q_{\mathrm{i}}$ is the internal quality factor, and $Q_{\mathrm{c}}$ is the coupling quality factor. Since the pump separation scheme relies on placing the pump tone near bottom of the resonance sweep, this shallowing of the resonance depth will place a limit on degree of pump separation.
@@
We then operated the same resonator as an amplifier, choosing the operating point using the pump separation scheme described earlier. The top subfigure of figure 5 shows the resonance sweep with the same power as the operating point using the pump tone as the sweeping tone; the bottom subfigure shows the amplification profile of the signal tone with the pump tone fixed at the operating point, which is indicated by the sharp spike at the centre of the figure. The operating point was given by using a pump power of $-18.00\,\mathrm{dBm}$ measured just outside of the cryostat, and a pump frequency of $6.20525\,\mathrm{GHz}$. Far from resonance, the full power of the pump was reflected, giving a reflection coefficient of $-27\,\mathrm{dB}$; at the operating point, most of the pump power is transmitted and a decreased reflection coefficient of $-32.5\,\mathrm{dB}$ was obtained instead. Far from resonance, the resonator reflects all the incident power, and the reflection coefficient is determined mainly by coaxial cable loss in the cryostat. This value can thus be used as a normalisation factor to determine the amount of pump power reflected from the resonator relative to its input, in this case $-32.5\,\mathrm{dB}+27\,\mathrm{dB}=-5.5\,\mathrm{dB}$. As expected from the design of the pump separation scheme, the signal tone was amplified strongly around the pump tone with a maximum gain of $\sim 19\,\mathrm{dB}$ while the reflected pump tone was reduced by $5.5\,\mathrm{dB}$ due to its placement near the bottom of the resonance in reflection. Our pump separation scheme turned out to be highly successful. Nevertheless, we note that the degree of separation is limited by the depth of the resonance, which is in turn limited by the relative magnitudes of the coupling and intrinsic quality factors as discussed above.
III.2 Optimised Qc measurements

### v3

@@ I Introduction
Signal amplification in superconducting parametric amplifiers is achieved through wave-mixing processes induced by the underlying nonlinearity, both reactive and dissipative, of the superconducting thin-film devices[1, 8, 9]. At present, there are two main types of superconducting parametric amplifier technologies: Josephson Parametric Amplifiers which rely on the nonlinear Josephson junction inductance [10, 11, 12], and Kinetic Inductance Parametric Amplifiers which rely on the nonlinear kinetic inductance of superconducting thin-films [1, 13]. In this study, we focus on the latter. Experimentally, kinetic inductance amplifiers have demonstrated gains of $10-30\,\mathrm{dB}$ whilst achieving noise performance close to the standard quantum limit for linear amplifiers of half a quantum [14, 1, 15, 16]. By operating in degenerate mixing mode, these amplifiers have also been used to realise squeezed amplification and generate squeezed vacuum states [17]. Superconducting parametric amplifiers have been realised in both travelling-wave and resonator geometries. In general, amplifiers based on travelling-wave designs have wider bandwidths, whereas amplifiers based on resonator designs have lower pump power requirements and are less prone to lithographic defects such as shorts or breaks on long ($>0.5\,\mathrm{m}$) transmission lines with micron-scale features [1, 18, 2]. Regardless of geometry, central to the process of parametric amplification is the use of a strong pump tone: the nonlinear wave-mixing processes typically transfer energy from the strong pump tone to the weak signal tone in order to achieve amplification. It is important that this strong pump tone (as large as $-8.0\,\mathrm{dBm}$ at the input of the amplifiers in some instances [1]) is separated from the amplified signal tone so that no unwanted saturation or nonlinearity occurs in the downstream electronics, especially cryogenic high electron mobility transistor (HEMT) amplifiers. Further, in the case of four-wave-mixing, the pump frequency is located close to the peak of the signal gain profile. As a result, this region of maximum gain may be unusable due to contamination of the signal by the phase-noise of the pump.
@@ II Theory
Figure 1: Top subfigure: schematic diagram of the pump separation scheme. The pump tone and the signal tone are injected into port 1 of the circulator; the input tones are sent to the half-wave resonator amplifier which is connected to port 2 of the circulator; the transmission mode of the resonator amplifier, which contain both the amplified signal tone and the pump tone, are terminated at a matched $50\,\mathrm{\Omega}$ termination; the reflection mode of the resonator amplifier, which contains mostly the amplified signal tone, is received via port 3 of the circulator as output. Bottom subfigure: an example of the half-wave resonators measured in this study [9]. The resonator comprises a length of coplanar waveguide capacitively coupled in series between two contact pads. The length over which the coupling pad overlaps with the resonator line determines the coupling quality factor of this resonator.
@@
Figure 2: Squared S-parameter of the reflection mode of a half-wave two port resonator at different sweep powers, simulated using the theory presented in [22]. The sweep direction is from low frequency to high frequency, and the power increases from red to magenta. The curves are spaced linearly in the amplitude of the sweep tone such that the red curve at $5\,\mathrm{GHz}$ corresponds to negligibly small sweep power, and the cyan curve at $4.995\,\mathrm{GHz}$ corresponds to just enough power to produce a bifurcation discontinuity that is as deep as the resonance depth. Accordingly, the resonance response shows increasing distortion due to a reactive feedback process. Discontinuity in the resonance response develops when the sweep power is greater than that required to induce bifurcation.
The nonlinear response of thin-film resonators as parametric amplifiers has been well-described by recent research and can be understood in two conceptual steps: The first step is to understand the nonlinear response of the resonator driven by a strong pump tone. This problem is relevant to other applications of superconducting resonators and is discussed in detail in [23, 24], along with the physical origin of the nonlinearities. The second step is to understand the subsequent behaviour when additionally a weak signal is applied, as detailed in [22]. Equivalently, we can view the problem as one of establishing the operating point for the device by the pump, followed by the small signal behaviour around that point.
In what follows, we will apply the analysis presented in [22] to the superconducting resonator amplifier described in this paper. This analysis assumes the resonator amplifier can be modelled by a set of temporal coupled-mode equations, and that the resonant frequency and the Q-factor of the mode vary with the square of the mode amplitude (i.e. the energy stored in the resonator). The result is a generalised Duffing-type model of nonlinear resonator which takes into account the effect of reactive, dissipative, and rate-limited nonlinearities in the wave-mixing process. Using the model, we have simulated the behaviour of a resonator with total quality factor $Q_{\mathrm{r}}$ of $1000$, coupling quality factor $Q_{\mathrm{c}}$ of $1100$, and resonance frequency of $5\,\mathrm{GHz}$, which is representative of the devices fabricated in our laboratory. We have also assumed that the nonlinearity is purely reactive, i.e. only the resonant frequency changes with power; this is the ideal case, and the aim of this section is to present qualitatively the characteristic behaviour pertaining to parametric amplification and pump-signal separation. Both figures in this section are modelling S-parameter of the reflection mode of the resonator (received at port 3 of the circulator configured according to figure 1).
In the presence of a strong tone, the stored and dissipated energies modify the circuit parameters of the resonator, which in turn change the frequency of the underlying, not directly observed, resonance. This reactive process results in key features in the nonlinear response such as shifts in resonance frequencies, distortions in S-parameter profiles, and hysteresis and bifurcations: i.e. the resonator responds differently depending on whether the frequency is swept up or swept down. As shown in figure 2, as the sweep power is increased, from the red line to the magenta line, the resonance shifts and distorts largely as a result of reactive parametric feedback. When the applied power is greater than that required for bifurcation, a discontinuity appears in the observed transmission profile. As described by our recent work on resonator parametric amplifiers [22], the onset of this discontinuity is associated with the formation of operating points that can be used for high-gain, low-noise amplification. In practice, the optimum pump power is approximately the applied power at bifurcation, and the optimum pump frequency is approximately the frequency at which the discontinuity first appears.
@@
The parametric amplification of the signal tone is a result of the wave-mixing between the signal and the pump facilitated by the underlying nonlinearity. At high gain, in a resonator, this process generates amplified signals of similar magnitude propagating in both the transmission and the reflection directions. The top subfigure of figure 3 shows the nonlinear resonance response in reflection where the power of the sweeping tone is equal to the pump power of the amplifier’s operating point; the bottom subfigure shows the gain profile of the reflected signal tone as a function of frequency. The peak gain in the bottom subfigure is $17\,\mathrm{dB}$. The dashed line indicates the best operating frequency of the pump tone; in the top subfigure, it intersects the pump response at a frequency just above the discontinuity with a reflection coefficient of $-10.5\,\mathrm{dB}$. This simulation illustrates that by operating a half-wavelength reflection amplifier with pump frequency just above the bifurcation frequency, the pump power reflected is significantly weaker than the pump power incident on the device; the signal tone, however, is strongly amplified in both the transmission mode and the reflection mode.
III Devices and measurements
We have designed, fabricated, and packaged half-wave NbN resonators to experimentally realise the pump separation scheme. The resonators were fabricated in batches of related designs on $50.8\,\,\mathrm{mm}$ diameter, $225\,\,\mathrm{\mu m}$ thick, CZ silicon wafers. The NbN device layer was patterned via a lift-off process with image-reversal photoresist. These were then diced these into individual resonator chips after processing. The NbN films were deposited by reactive DC sputtering a niobium target in the presence of argon as the buffer gas and nitrogen as the reactive gas. All depositions were carried out in an ultra-high vacuum deposition system which has a base pressure of $2\times 10^{-10}\,\,\mathrm{Torr}$ or below. The measured superconducting transition temperatures of the films were $\sim 10\,\mathrm{K}$.
All devices tested in this study were half-wave resonators based on co-planar waveguide (CPW) geometries, as shown in the bottom subfigure of figure 1. The CPW conductor width was $5\,\mathrm{\mu m}$, and the gap width from conductor to ground was $50\,\mathrm{\mu m}$. These resonators were end coupled through capacitive small re-entrant lines. The film thicknesses were controlled by calibrating and varying the deposition time. All devices in this study had thicknesses of $\sim 100\,\mathrm{nm}$.
For testing, the resonators were bonded using aluminium wires to a gold-plated copper enclosure with SMA connectors. The enclosure was designed such that its resonance frequencies do not overlap significantly with the resonance frequencies of the half-wave resonators [25]. The packaged amplifier and the cryogenic circulator were arranged according to the top subfigure of figure 1. The cryogenic circulator used here was a QuinStar QCY series single-junction circulator with Mu-metal shield. Both the amplifier and the circulator were attached to the cold stage of an adiabatic demagnetization refrigerator. Stainless steel coaxial cables were used in the cryogenic system, thermally anchored at $35\,\mathrm{K}$, $4\,\mathrm{K}$, $1\,\mathrm{K}$, and $50\,\mathrm{mK}$ stages [9]. HEMT amplifiers were not used in the signal chain since added-noise temperature measurement was not an aim of this study. The pump tone was generated and measured using a vector network analyser (VNA) and the signal tone was generated and measured using another VNA. The two VNAs were phase-locked to a common $10\,\mathrm{MHz}$ reference tone. The signal and the pump tone were combined at the input of the cryostat using a multiband two-way power combiner. In all measurements, both the pump and the signal tones were injected via port 1 and measured via port 3 of the circulator, where the ports were configured according to figure 1. Unless otherwise specified, all measurements were taken at cryogenic temperature of $0.1\,\mathrm{K}$, maintained using an adiabatic demagnetization refrigerator with a pulse tube cooler.
III.1 Amplifier operating point
@@
Once the hardware components are configured properly, the experimenter still needs to find the operating point of the superconducting resonator parametric amplifier in terms of pump frequency and power to achieve both high gain and high degree of pump separation. Finding the operating point appears, at first sight, to be a complicated task. As a strong pump is swept across the notional resonance frequency, the energies stored and dissipated in the resonator change. Although the underlying resonance simply shifts and broadens, retaining a simple Lorentzian form, this is not what is observed by the experimenter in a swept measurement, which exhibits distortion and bifurcation. This effect is highly reproducible, and well described by our theory [22], and so we have been able to establish a set-up procedure, which has proven to be exceedingly dependable and reproducible when applied to a wide range of resonator configurations and materials. The ruggedness of the operation is one of the key attractions of the resonator approach. The experimental procedure is as follows:
@@
III.2 Proof-of-concept measurements
@@
In our first set of experiments, we applied the pump separation scheme to a half-wave NbN resonator with quality factor of $Q_{\mathrm{r}}\sim 30000$ (when measured using a weak sweep tone). Figure 4 shows measurements of S-parameter where the sweep tone was injected via port 1 of the circulator and measured via port 3 of the circulator. The ports were configured according to figure 1. The measured values were offset by a $20\,\mathrm{dB}$ attenuator before the input of the cryostat, coaxial cable loss in the cryostat, and circulator loss at the cold stage of the cryostat. (This is the only measurement where there is an attenuator in the pump path. The attenuator was included here to enable measurement of the resonance at very low sweep powers.) The measured S-parameter resonance sweeps display additional features compared to the idealised kinetic inductance nonlinear behaviour shown in figure 2. Firstly, the depth of the resonance decreases with increasing power. This indicates increase in dissipation with readout power, which may be explained by the non-equilibrium quasiparticle heating mechanism [26, 27, 8]. Additionally, the shift in resonance frequency was initially in the positive frequency direction and then changed back to the negative frequency direction as the power was increased further. This complicated effect has previously been observed and discussed in [9] in the context of Ti, Nb, and NbN resonators. Overall, at high powers, the nonlinearity is dominated by the kinetic inductance nonlinearity and displays its key features such as the bifurcation discontinuity. As seen in figure 4, when the sweep power increases to $2\,\mathrm{dBm}$ and above, the S-parameter shows a discontinuity, which is the reactive bifurcation.
The fact that the depth of resonance becomes shallower as the sweep power increases indicates a decrease in the internal quality factor of the resonator as a result of increased dissipation. From well-established resonator theories [24, 23], the depth of the resonance is given by
$|S|^{2}$ $=\left(1-\frac{Q_{\mathrm{r}}}{Q_{\mathrm{c}}}\right)^{2}\,,$ (1)
where $Q_{\mathrm{r}}$ is the total quality factor given by $Q_{\mathrm{r}}^{-1}=Q_{\mathrm{i}}^{-1}+Q_{\mathrm{c}}^{-1}$, $Q_{\mathrm{i}}$ is the internal quality factor, and $Q_{\mathrm{c}}$ is the coupling quality factor. Since the pump separation scheme relies on placing the pump tone near the bottom of the resonance sweep, this shallowing of the resonance depth will place a limit on degree of pump separation.
@@
We then operated the same resonator as an amplifier, choosing the operating point using the pump separation scheme described earlier. The top subfigure of figure 5 shows frequency sweep of the resonance using the pump tone as the sweeping tone. The pump tone has the same power as the amplifier operating point and is the only tone present in this resonance sweep. The bottom subfigure shows the amplification profile of the signal tone with the pump tone fixed at the operating point. The sharp spike at the centre of the bottom subfigure indicates the frequency location of the strong pump tone and does not represent actual gain. The operating point was given by using a pump power of $-18.00\,\mathrm{dBm}$ measured just outside of the cryostat, and a pump frequency of $6.20525\,\mathrm{GHz}$. Far from resonance, the full power of the pump was reflected, giving a reflection coefficient of $-27\,\mathrm{dB}$; at the operating point, most of the pump power is transmitted and a decreased reflection coefficient of $-32.5\,\mathrm{dB}$ was obtained instead. The same reflection coefficient was obtained both by reading off the resonance sweep at the frequency of the operating point (top subfigure) as well as by measuring the pump S-parameters during an amplification sweep with the pumping VNA set to continuous wave mode (bottom subfigure). Far from resonance, the resonator reflects all the incident power, and the reflection coefficient is determined mainly by coaxial cable loss in the cryostat. This value can therefore be used as a normalisation factor to determine the amount of pump power reflected from the resonator relative to its input, in this case $-32.5\,\mathrm{dB}+27\,\mathrm{dB}=-5.5\,\mathrm{dB}$. As expected from the design of the pump separation scheme, the signal tone was amplified strongly around the pump tone with a maximum gain of $\sim 19\,\mathrm{dB}$ while the reflected pump tone was reduced by $5.5\,\mathrm{dB}$ due to its placement near the bottom of the resonance in reflection. Our pump separation scheme turned out to be highly successful. Nevertheless, we note that the degree of separation is limited by the depth of the resonance, which is in turn limited by the relative magnitudes of the coupling and intrinsic quality factors as discussed above.
III.3 Optimised Qc measurements